Ok so I've been very frustrated trying to solve this problem, Please help. Test is tomorrow.
Trimming a Garage Door: A carpenter used 30Ft of molding in three pieces to trim a garage door. If the long piece was 2ft longer than twice the length of each shorter piece, then how long was each piece?
(only three sides to this trim. Two vertical and one long horizontal piece.)
I came up with an answer but it doesn't work. My answer was 8ft 8ft and 14ft. but the problem states that each short piece has to be twice the length of the long piece plus 2ft more.
Here is what i did:
2ft + 2L = 30ft solve and i got 14ft for the long and then 8ft for each short piece.
My book gives the answer which comes out to be 7ft 7ft and 16ft but I just cant figure out how to get that on my own.
Thanks for reading, help is much appreciated.
-Jeremy

Answer by scott8148(6628)

s = short length

"the long piece was 2ft longer than twice the length of each shorter piece" ___ L = 2s + 2

"A carpenter used 30Ft of molding in three pieces" ___ s + s + L = 30 ___ s + s + (2s + 2) = 30

4s + 2 = 30 ___ s = 7

L = 2(7) + 2 = 16